Solve the system by addition
-3x+4y=-16
x-2y=12

Answer by funmath(2873)

Solve the system by addition
-3x+4y=-16 R1
x-2y=12 R2
Multiply R2 by 3 and add it to R1 and the x's will be eliminated so you can solve for y.
3R2===> 3(x-2y)=3(12)===> 3x-6y=36
R1+3R2 is
-3x+4y=-16
3x-6y=36
----------
0-2y=20
-2y=20
-2y/-2=20/-2
y=-10
Now substitute 10 in for y in R2 (because it's easier than R1) and solve for x:
x-2(-10)=12
x+20=12
x+20-20=12-20
x=-8
The solution to the system is (x,y)=(-8,-10)
Happy Calculating!!!!